OpenStudy (xximhisgirlxx):

Make a conjecture about the next item in the sequence. 3, 6, 12, 24, … 32 36 48 72

OpenStudy (studygurl14):

What is the change that gets from 3 to 6? What is the change that gets from 6 to 12? If you see a pattern, just apply that pattern rule to 24 to get the next number

OpenStudy (ayyookyndall):

Every time it is x2

OpenStudy (studygurl14):

6/3 = 2 12/6 = 2 24/12 = 2 So...24 • 2 = 48 C

OpenStudy (xximhisgirlxx):

Is it 48?

OpenStudy (ayyookyndall):

Yes.

OpenStudy (xximhisgirlxx):

Y'all mind helping with a few more, please?

OpenStudy (ayyookyndall):

Nope.

OpenStudy (ayyookyndall):

Ask away